a plane flies 1440 miles with a tail wind in 4 hours. The plane makes the return trip against the same wind and takes 4 1/2 hours. find the speed of the plane and of the wind.
Answer by mananth(16949) 
 Plane speed =x mph
wind speed =y mph
against wind 4.5 hours
with wind 4 hours
Distance against 1440 miles distance with 1440 miles
t=d/r against wind -
1,440.00 / ( x - y )= 4.50
4.50 ( x - y ) = 1,440.00
4.50 x - 4.50 y = 1,440.00 ....................1
1,440.00 / ( x + y )= 4.00
4.00 ( x + y ) = 1,440.00
4.00 x + 4.00 y = 1,440.00 ...............2
Multiply (1) by 4.00
Multiply (2) by 4.50
we get
18.00 x + -18.00 y = 5,760.00
18.00 x + 18.00 y = 6,480.00
36.00 x = 12,240.00
/ 36.00
x = 340.00 mph
plug value of x in (1) y
4.50 x -4.50 y = 1,440.00
1,530.00 -4.50 -1,530.00 = 1,440.00
-4.50 y = 1,440.00
-4.50 y = -90.00 mph
y = 20.00
Plane speed 340.00 mph
wind speed 20.00 mph
m.ananth@hotmail.ca
|
|
|
| |